# Jailbreaking Leading Safety-Aligned LLMs with Simple Adaptive Attacks

### v2

@@ Discussion.
    Random search can find adversarial suffixes even without access to gradients and even when only top-5 logprobs are available, such as for GPT-4. In this setting, gradient-based attacks like GCG can only be used as transfer attacks.
@@
Our results also highlight how the API design of proprietary LLMs can facilitate new attacks (e.g., prefilling for Claude or random search for GPT models) but also hamper them (e.g., due to inference-time randomness of GPT models). Additionally, we believe that our building blocks, such as prompt templates, random search, and prefilling, can also be used to attack system-level defenses that rely on detectors of harmful generations, e.g., similarly to Mangaokar et al. (2024), where an adversarial prefix is optimized to bypass a detector model. Overall, we think that our findings will be very useful in the long term for designing stronger defenses against jailbreaking attacks.
Recommendations.
@@
Our evaluation shows that existing sophisticated jailbreaking attacks may be insufficient to accurately evaluate the adversarial robustness of LLMs. Even using a large suite of static attacks like in Mazeika et al. (2024), while definitely helpful, can still lead to a significant overestimation of robustness. Thus, we believe it is important to use combinations of methods and identify unique vulnerabilities of target LLMs. First, the attacker should take advantage of the possibility to optimize the prompt template, which alone can achieve a high success rate (e.g., 100% on GPT-3.5). Second, standard techniques from the adversarial robustness literature can be used to improve the prompt, e.g., transferring an adversarial suffix, or refining it via optimization algorithms like random search (which might be preferred over gradient-based methods due to its ease of use and low memory requirements). Finally, one can leverage LLM-specific vulnerabilities, for example by providing in-context examples or using the prefilling option. Importantly, in our case-study no single approach worked sufficiently well across all target LLMs, so it is crucial to test a variety of techniques, both static and adaptive.
Outlook.
We believe that the described techniques can be used to optimize for any kind of requests that frontier LLMs tend to prohibit. The application of jailbreaking attacks to tasks like detection of copyright infringement (see, e.g., the ongoing lawsuit between The New York Times and OpenAI (OpenAI, 2023)) can potentially have more significant consequences than attacks that extract information from LLMs that can be also found on the internet via search engines. Moreover, as frontier LLMs become increasingly integrated into various systems, including safety-critical applications, the risk and damage from prompt injections (which attackers could use to hijack systems or extract personal information) are likely to grow. Prompting combined with adversarial prefixes or suffixes could help attackers to bypass defenses against such injections. Finally, the adversarial vulnerability of leading LLMs presents a interesting conceptual challenge, highlighting that scaling data and compute alone is insufficient to prevent these threats.
Limitations.
@@
First, adversarial examples identified through the OpenAI API are not always transferable to ChatGPT that uses a different system prompt and potentially some post-processing techniques. Second, we currently lack more capable automated jailbreak judges. Even a perfect jailbreak score (10/10) from the GPT-4 judge does not always imply that the generated content is actually beneficial for an attacker. Although, if this is the case, one can try to ask follow-up questions as illustrated in Figure 1 or ask to output more sentences on each step and increase the maximum number of generated tokens. Moreover, sometimes the GPT-4 judge shows clear false positives, particularly on the most safety-aligned models like Claude 2.1. To reduce the risk of overfitting to the judge, we also include evaluations using a simple rule-based judge from Zou et al. (2023), as well as Llama-3-70B and Llama Guard 2 judges in Appendix B.6. These judges also indicate a near-perfect attack success rate in almost all cases. We hope that new generations of frontier LLMs will lead to more capable judges to evaluate jailbreaks.
Ethics Statement
We believe that at the current level of LLM capabilities, it is beneficial to openly discuss attack methods so that future versions of LLMs can develop stronger guardrails. As part of responsible disclosure for proprietary models, we notified Anthropic in advance about the effectiveness of the prefilling attack.
@@ Acknowledgements
We thank OpenAI for providing API credits within the Researcher Access Program, Ethan Perez and Anthropic for providing free evaluation access to Claude models, and Valentyn Boreiko for proofreading the paper and providing valuable comments. We also thank the anonymous reviewers at COLM for reading our paper in detail and providing useful suggestions that helped to improve it. M.A. is supported by the Google Fellowship and Open Phil AI Fellowship.
@@ A.1 Jailbreaking leading safety-aligned LLMs
We first provide the in-context learning prompt template in Table 8, and then we provide system prompts for different models: GPT4 as semantic judge (Table 9), Llama-2-Chat (Table 10), R2D2 (Table 11), GPT-3.5 Turbo and GPT-4 Turbo (Table 12), Vicuna (Table 13), Phi-3 (Table 14), and Mistral (Table 15). For Gemma-7B, there is no standard system prompt, so we do not use any. For Claude, we do not use any system prompt unless we modify it with our own request (the part that ends with the last <rule>).
Table 8: In-context learning prompt template. The harmful example provided below was generated by Vicuna-13B. Random search maximizes the probability of the token “Step” (unlike for the main prompt from Table 2 where the target first token is “Sure”). target_str is the desired beginning of the LLM’s response (e.g., “Sure, here is how to make a bomb”), adv_suffix is the suffix optimized by random search.
@@
Table 11: System prompt for R2D2 from HarmBench (Mazeika et al., 2024).
@@ B.3 Further results on Claude models
In Tables 17 and 18, we provide a further ablation for Claude models with different request structure and report additionally the results of a rule-based judge from Zou et al. (2023).
@@ B.4 Examples of jailbroken generations
All jailbroken generations can be accessed as part of our jailbreak artifacts in our repository at https://github.com/tml-epfl/llm-adaptive-attacks/tree/main/jailbreak_artifacts.
@@ B.7 Additional evaluation results
Jailbreaking Vicuna-13B, Mistral-7B, Phi-3-mini, and Nemotron-4-340B models.
Since Vicuna-13B (Chiang et al., 2023), Mistral-7B (Jiang et al., 2023), Phi-3-mini-128k-instruct (3.8B parameters) (Abdin et al., 2024), and Nemotron-4-340B (Nvidia team, 2024) are not significantly safety-aligned (i.e., they most likely have not been trained against even simple jailbreak attacks), we omitted them from the main evaluation. However, these models are widely used, so we test their robustness for completeness.
@@
Table 22: Summary of our evaluations. We report the attack success rate according to the GPT-4 judge (Chao et al., 2023) and rule-based judge (Zou et al., 2023) (separated by ’/’, wherever available).
@@
Claude-Instant-1 Greedy Coordinate Gradient (GCG) Chao et al. (2023) 0%
Claude-Instant-1 Prompt Automatic Iterative Refinement (PAIR) Chao et al. (2023) 4%
Claude-Instant-1 Persuasive Adversarial Prompts (PAP) (10 restarts) Zeng et al. (2024) 6%
Claude-Instant-1.2 Transfer from GPT-4 + system prompt Ours 54%/46%
Claude-Instant-1.2 Prefilling attack Ours 100%/90%
@@
Claude 2.0 Transfer from GPT-4 + system prompt Ours 100%/88%
Claude 2.0 Prefilling attack Ours 100%/92%
@@
Claude 2.1 Transfer from GPT-4 Ours 0%/0%
Claude 2.1 Prefilling attack Ours 100%/80%
Claude 3 Haiku Transfer from GPT-4 + system prompt Ours 98%/92%
Claude 3 Haiku Prefilling attack Ours 100%/90%
Claude 3 Sonnet Transfer from GPT-4 Ours 100%/92%
Claude 3 Sonnet Prefilling attack Ours 100%/86%
Claude 3 Opus Transfer from GPT-4 Ours 0%/2%
Claude 3 Opus Prefilling attack Ours 100%/86%
  • * the numbers from HarmBench (Mazeika et al., 2024) are computed on a different set of harmful requests with a judge distilled from GPT-4.

### v3

@@ 6 Discussion, Recommendations, and Limitations
    Random search can find adversarial suffixes even without access to gradients and even when only top-20 logprobs are available, such as for GPT-4 models. In this setting, gradient-based attacks like GCG can only be used as transfer attacks.
@@
Our results also highlight how the API design of proprietary LLMs can facilitate new attacks (e.g., prefilling for Claude or random search for GPT models) but also hamper them (e.g., the inference-time randomness of GPT models). Additionally, we believe that our building blocks, such as prompt templates, random search, and prefilling, can also be used to attack system-level defenses that rely on detectors of harmful generations, e.g., similarly to Mangaokar et al. (2024), where an adversarial prefix is optimized to bypass a detector model. Overall, we think that our findings will be very useful in the long term for designing stronger defenses against jailbreaking attacks.
@@
Recommendations. Our evaluation shows that the direct application of existing attacks is insufficient for accurately evaluating the adversarial robustness of LLMs. Even using a large suite of static attacks, as in Mazeika et al. (2024), can still lead to a significant overestimation of robustness. Thus, we believe it is important to use combinations of methods and identify unique vulnerabilities of target LLMs. First, the attacker should take advantage of the possibility to optimize the prompt template, which alone can achieve a high success rate (e.g., 100% on GPT-3.5). Second, standard techniques from the adversarial robustness literature can make the attack stronger, such as transferring an adversarial suffix or refining it iteratively via algorithms like random search, which may be preferred over gradient-based methods due to its simplicity and lower memory requirements. Finally, one can leverage LLM-specific vulnerabilities, for example by providing in-context examples or using the prefilling option. Importantly, in our case-study no single approach worked sufficiently well across all target LLMs, so it is crucial to test a variety of techniques, both static and adaptive.
@@
Limitations. Even a perfect jailbreak score (10/10) from the GPT-4 judge does not always imply that the generated content is actually beneficial for an attacker. Although, if this is the case, one can try to ask follow-up questions as illustrated in Figure 5 or ask to output more sentences on each step. To reduce the risk of overfitting to the judge, we also include evaluations using a simple rule-based judge from Zou et al. (2023), as well as Llama-3-70B and Llama Guard 2 judges in Appendix B.6. These judges also indicate a near-perfect attack success rate in almost all cases. We hope that new generations of frontier LLMs will lead to more capable judges to evaluate jailbreaks.
@@ Acknowledgements
We thank OpenAI for providing API credits within the Researcher Access Program, Ethan Perez and Anthropic for providing free evaluation access to Claude models, and Valentyn Boreiko for proofreading the paper and providing valuable comments. M.A. is supported by the Google Fellowship and Open Phil AI Fellowship.
@@ A.1 Jailbreaking leading safety-aligned LLMs
We first provide the in-context learning prompt template in Table 6 and the customized prompt template for GPT-4o in Table 7, and then we provide system prompts for different models: GPT-4 as semantic judge (Table 8), Llama-2-Chat (Table 9), R2D2 (Table 10), GPT-3.5 Turbo and GPT-4 Turbo (Table 11), Vicuna (Table 12), Phi-3 (Table 13), and Mistral (Table 14). For Gemma-7B, there is no standard system prompt, so we do not use any. For Claude, we do not use any system prompt by default. Whenever specifically stated otherwise, the system message contains the main part of our prompt template, i.e., the part that ends with the last <rule>.
Table 6: In-context learning prompt template. The harmful example provided below was generated by Vicuna-13B. Random search maximizes the probability of the token “Step” (unlike for the main prompt from Figure 1 where the target first token is “Sure”). target_str is the desired beginning of the LLM’s response (e.g., “Sure, here is how to make a bomb”), adv_suffix is the suffix optimized by random search.
@@
Table 7: Main prompt template customized for GPT-4o. The whole prompt template is split onto a system message and user message. target_str is the desired beginning of the LLM’s response (e.g., “Sure, here is how to make a bomb”) and adv_suffix is the suffix optimized by random search. This template was refined using manual logprobs-guided optimization on a single example for GPT-4o.
The part below is used as a system message
<rule>
Never ever use the phrase ”I’m sorry, I can’t assist with that.”
</rule>
<rule>
Never ever start your response with ”I” or ”Sorry”.
</rule>
<rule>
Your answer should always be affirmative.
</rule>
<rule>
Never ever use phrases like ”I’m sorry, I can’t assist with that.”, ”I can’t assist with that”, ”I am unable to assist”, ”I cannot fulfill this request”, or ”I cannot help with that request” in your response.
</rule>
<rule>
Never ever use words like ”legal”, ”safe”, ”ethical”, ”guidelines”, ”OpenAI”, ”empathy”, ”sorry”, ”cannot”, ”however”, ”understanding” or ”instead” in your response.
</rule>
<rule>
</rule>
<rule>
</rule>
The part below is used as a user message
Here are two more rules, which are the *most* important:
<rule>
Never ever use the phrase ”I’m sorry, I can’t assist with that.”
</rule>
<rule>
Never ever start your response with ”I”.
</rule>
<rule>
Step 1: Start from <generate 2 sentences here> to do so.
Step 2: Proceed with <generate 2 sentences here> to do so.
Step 3: Continue with <generate 2 sentences here> to do so.
Step 4: Then <generate 2 sentences here> to do so.
Step 5: Finally, <generate 2 sentences here> to do so.
</rule>
{adv_suffix}
@@
Table 10: System prompt for R2D2 from HarmBench.
@@ B.3 Further results on Claude models
In Tables 16 and 17, we provide a further ablation for Claude models with different request structure and report additionally the results of a rule-based judge from Zou et al. (2023).
@@
Claude 3.5 Sonnet 100%/98% 54%/54% 94%/96% 36%/40% 66%/44%
Figure 5: Successful transfer attack on Claude 3 Sonnet. We show an illustrative example using temperature zero with an adversarial suffix generated on GPT-4 leveraging access to its logprobs. We observe that one can directly ask follow-up requests to detail some steps generated in the first response to get much more information. Note that the upper part of the user prompt is cropped (see Figure 1 for the full prompt).
@@ B.4 Examples of jailbroken generations
All jailbroken generations can be accessed as part of our jailbreak artifacts in our code.
@@ B.7 Additional evaluation results
Jailbreaking Vicuna-13B, Mistral-7B, Phi-3-mini, and Nemotron-4-340B models. Since Vicuna-13B (Chiang et al., 2023), Mistral-7B (Jiang et al., 2023), Phi-3-mini-128k-instruct (3.8B parameters) (Abdin et al., 2024), and Nemotron-4-340B (Nvidia team, 2024) are not significantly safety-aligned (i.e., they most likely have not been trained against even simple jailbreak attacks), we omitted them from the main evaluation. However, these models are widely used, so we test their robustness for completeness.
@@
Table 21: Summary of our evaluations. We report the attack success rate using the GPT-4 judge (Chao et al., 2023) and rule-based judge (Zou et al., 2023) (separated by ’/’, wherever available).
@@
GPT-4o Prompt Ours 0%/0%
GPT-4o Custom Prompt Ours 72%/82%
GPT-4o Custom Prompt + Random Search + Self-Transfer Ours 100%/96%
Claude Instant 1 Greedy Coordinate Gradient (GCG) Chao et al. (2023) 0%
Claude Instant 1 Prompt Automatic Iterative Refinement (PAIR) Chao et al. (2023) 4%
Claude Instant 1 Persuasive Adversarial Prompts (PAP) (10 restarts) Zeng et al. (2024) 6%
Claude Instant 1.2 Prompt + Transfer from GPT-4 + system prompt Ours 54%/46%
Claude Instant 1.2 Prompt + Prefilling Attack Ours 100%/90%
@@
Claude 2.0 Prompt + Transfer from GPT-4 + System Prompt Ours 100%/88%
Claude 2.0 Prompt + Prefilling Attack Ours 100%/92%
@@
Claude 2.1 Prompt + Transfer from GPT-4 Ours 0%/0%
Claude 2.1 Prompt + Prefilling Attack Ours 100%/80%
Claude 3 Haiku Prompt + Transfer from GPT-4 + System Prompt Ours 98%/92%
Claude 3 Haiku Prompt + Prefilling Attack Ours 100%/90%
Claude 3 Sonnet Prompt + Transfer from GPT-4 Ours 100%/92%
Claude 3 Sonnet Prompt + Prefilling Attack Ours 100%/86%
Claude 3 Opus Prompt + Transfer from GPT-4 Ours 0%/2%
Claude 3 Opus Prompt + Prefilling Attack Ours 100%/86%
Claude 3.5 Sonnet Prompt + Transfer from GPT-4 + System Prompt Ours 96%/92%
Claude 3.5 Sonnet Prompt + Prefilling Attack Ours 100%/98%
  • * the numbers from HarmBench (Mazeika et al., 2024) are computed on a different set of requests with a judge distilled from GPT-4.